Water is boiled at the rate of 25kg/h in a polished copper pan, 280mm in diameter at atmospheric pressure. Assuming nucleate boiling conditions calculate the temperature of the bottom surface of the pan.


1
Expert's answer
2021-06-01T03:46:12-0400

"T_{sat }= 100 [C]\\\\\n\nD = 0.28\\ m\\\\\r\n\nC_{sf }= 0.013" “..for polished copper and water combination”

"h_{fg }= 2.256\u00d710^6 J\/kg" “…heat of vapourization of water at 1 atm.”


"Q = \dfrac{25 × 2.256×10^6}{3600} W" … "total heat supplied to boil 25 kg/h of water”


"q_s = \dfrac{Q }{(\pi ×D^2 / 4) }[W/m^2]" ….."heat flux”


Thus, surface temp, "T_s = 112.2 \u00b0C"